# On optimal solutions of classical and sliced Wasserstein GANs with non-Gaussian data

### v1

@@ III Main Results
In Sec. III-A, we analytically solve WGAN (1) with the non-linear generator and non-Gaussian data under $d=1$ when $q=2$. With large $d$ in Sec. III-B, we show that the linear generators can asymptotically optimal for both original and unprojected sliced WGANs (II) and (6) by solving corresponding upper-bounds.
III-A Results for WGAN under one dimension
@@
Proof:
The solution of the first sub-problem is indeed (9) while that for second sub-problem is (10). The condition (8) is obtained by comparing the values of the two subproblems. The following proof is given in Sec. IV-A. ∎
@@
where $\Phi$ is the CDF of Gaussian $N(0,1)$. However, now $\Psi$ has a jump from 0 to $1/2$ at $v=0$ and does not meet the setting of Theorem 1 since it is neither continuous nor strictly increasing. We need the following modification.
@@
Assume $\mu$ has the setting as in Theorem 1 and $\Psi$ is given by (12). If
$\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)$
$\geq$ $\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right),$ (13)
@@
The proof of Theorem 2 is in Appendix A. For the linear generator $h(Z)=Z$ as [5, eqn. (27)][8][9], one can simplify the results in Theorem 1 and also get alternative closed-form formula as follows.
@@
Assume $\mu$ has the setting as in Theorem 1 and consider the linear case $h(Z)=Z$ in (7). Then population WGAN (1) has a unique minimizer for $(\theta_{1},\theta_{2})\in\mathbb{R}\times\mathbb{R}$ as
@@
The proof of Corollary 1 is in Appendix B.
Here we briefly compare our proposed WGAN solutions with the results for the LQG setting in [8, Theorem 1]. Unlike [8], our results can deal with non-Gaussian data distribution. Moreover, we can recover the result in [8] when $q=2,d=1$. Indeed, if $X\sim\mathcal{N}(0,\sigma^{2})$, by looking into (16), we have $\Phi^{-1}(F_{\mu}(X))=\Phi^{-1}(\Phi(X/\sigma))=X/\sigma.$ Plugging this into our solution in (16) shows that $\theta_{1}^{*}=0$ and $\theta_{2}^{*}=\mathbb{E}[X\frac{X}{\sigma}]=\sigma$, coinciding with the result in [8] for $d=r=1$. We emphasize that neither of our and the results in [8] subsume the other as a special case as [8] considers general dimension $d$ while our work is not restricted to Gaussian data distribution. Since WGAN output is already Gaussian $\Theta Z\sim\mathcal{N}(0,\mathbf{K}^{\prime})$, it is shown in [8, Theorem 1] that for this benchmark, the optimal WGAN solution happens to be the $r$-PCA solution of Gaussian $X$. That is, optimal generator matrix $\Theta$ fulfills that $\mathbf{K}^{\prime}=\Theta\Theta^{T}$ is a rank $r$ matrix and $\mathbf{K}^{\prime}$ and $\mathbf{K}$ share the same largest $r$ eigenvalues and the corresponding eigenvectors.
Finally, we present our result for non-linear generators and first-order Wasserstein distance ($q=1$). We have
@@
Following the settings in Theorem 1, the minimizer $(\theta^{*}_{1},\theta^{*}_{2})$ of the population WGAN (1) with $q=1,d=1$ meet the following necessary conditions
@@
III-B Results for sliced WGAN under high dimension
We first focus on the asymptotic optimal generator of (II) as defined in Sec. II. Our main result shows that the linear generator can be asymptotically optimal for the original sliced WGAN (II) when $q=2$.
@@ III-B Results for sliced WGAN under high dimension
For the second-order sliced WGAN (II), if data $X$ satisfy the weak dependence condition in [16, Definition 1], the linear generator (3) is asymptotically optimal with the parameters
@@
For the proof sketch, rather than finding the closed-form solution of the original sliced WGAN (II) directly, we solve its upper-bounds with bounded-error and approximations to overcome the the complicated integration and projection of (II). First, by replacing general generator $G_{\Theta}(Z)$ with linear generator (3), (II) is upper-bounded by
$\min_{\Theta}\left(\int_{\omega\in\Omega}\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|^{q}]d\omega\right)^{\frac{1}{q}}.$ (20)
Motivated by [17] we replace the uniformly distributed $\omega\in\Omega$ in (20) with Gaussian projections $\omega\in\Omega_{G}$, that is, $\omega$ is generated according to a zero-mean Gaussian distribution with covariance matrix $(1/d)\mathbf{I}_{d}$ and aim at solving
$\min_{\Theta}\left(\int_{\omega\in\Omega_{G}}\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|^{q}]d\omega\right)^{\frac{1}{q}},$ (21)
@@
$\min_{\Theta}\left(\int_{\omega\in\Omega_{G}}\inf_{\pi}\mathbb{E}_{\pi}[|\hat{X}_{G}-\omega^{T}\Theta Z|^{q}]d\omega\right)^{1/q}$ (22)
is bounded by a function $f^{\mu}(d)=O(d^{-1/8})$, where the marginals of $\pi$ are $\mathcal{N}(0,\tilde{\sigma}^{2}_{x})$ and $\nu^{\Theta}_{\omega}$. Solving (22) with the linear generator (3), the optimal $\Theta$ equals to that of
@@
As summary to the current steps, the upper-bounds of (II) under $q=2$ are
$\min_{\Theta}\left(\int_{\omega\in\Omega}\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}G_{\Theta}(Z)|^{2}]d\omega\right)^{\frac{1}{2}}$
$\leq$ $\min_{\Theta}\left(\int_{\omega\in\Omega_{G}}\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}]d\omega\right)^{\frac{1}{2}}$
@@
where the optimal $\Theta$ in (III-B) can be solved by (24). Then, the final upper-bound in (III-B) becomes
@@
Indeed one can show that (26) tends to zero as $d\rightarrow\infty$, and thus linear generator in the second upper-bound of (III-B) is asymptomatically optimal for (II). Thus, in the sense of asymptotic optimality defined in Sec. II, even with the non-linear generator, the value of sliced WGAN under very high dimensional non-Gaussian data cannot be smaller than that from proposed linear generator (19).
@@
Then, we prove that the obejective of (27) is Schur-convex [19], so the global minimum of (27) is acheived when all the diagnol elements of $\mathbf{S}$ are equal. Thus, we find the optimal parameters (19) of linear generator, and obtain the upper bound value (26). The full proof of solving the upper bound is given in Sec. IV-B.
@@
From [16], under the weak dependence condition, the elements of data vector $X$ can follow Gaussian process, Bernoulli shift and autoregressive (AR) model. For channel prediction, AR model-based method is widely used by using past channel data without requiring current channel measure via pilot [20].
@@
As a sanity check, when $d=2$ and $q=2$, (22) with linear generator can also be solved by elliptic integrals. The solution are the same as (19). The proof is given in appendix E.
@@
Under assumption of Theorem 3 with $\mu$ vanishing on hypersurfaces in $\mathbb{R}^{d}$, for the unprojected sliced WGAN (6), the linear generator (3) with the same parameters as (19) is asymptotically optimal when $q=2$.
@@
We show that (6) is indeed equivalent to (II) when $q=2$ by the following lemma.
@@
For any $\omega\in\Omega$, the values of the inner-discriminator probelms of (6) and (II) with the linear generator are the same when $q=2$,
$\inf_{\pi\in\Pi(\mu,\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}]$ (28)
$=$ $\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}].$
We prove that the optimal transport function $\pi^{*}$ of the L.H.S. in (28) equals to that of R.H.S. in (28). First, the L.H.S. in (28) belongs to the general multi-to-one-dimensional optimal transport problems with closed-form solutions identified in [21]. Then, we identify its analytical properties to prove that the optimal transport function of the L.H.S. of (28) is well-defined and equals to that of the R.H.S.. The detailed proof of Lemma 2 is in Sec. IV-C. Finally, we can prove the desired result for (6) following steps to prove that for original sliced WGAN (II) in Theorem 3. ∎
For application, our upcoming simulations in Sec. V shows that the linear generator (19) can be a low-complexity generator with empirical data. For data without weak dependence condition, the state-of-art WGAN solver can be obtained via gradient-flow of stochastic differential equation [22][23], and our optimal can be a low-complexity initialization and has the potential for faster convergence.
@@ IV-A Proof for Theorem 1
In the following, we first solve the first sub-problem with $\theta_{2}\in\mathbb{R}_{+}$ in (11). With $d=1$, we recall the following result for the inner-discriminator problem (2) of (1), by taking $\nu^{\Theta}:=P_{G_{\Theta}(z)}$, the measure generated by the generator NN (with parameter $\Theta:=(\theta_{1},\theta_{2})$ in (7)). Let $F_{\mu}$ and $F_{\nu^{\Theta}}$ denote the cumulative distribution functions (CDFs) of the measures $\mu$ and $\nu^{\Theta}$ on $\mathbb{R}$.
@@
For the inner discriminator problem (2) with $q=2,d=1$, $\mathbb{E}_{\mu}[P(X,\theta_{1},\theta_{2})]$ for given $(\mu,\theta_{1},\theta_{2})$ equals to (30), where $t^{\Theta}(x)$ is defined as in (29). Now we need to find a closed-form $t^{\Theta}(x)$ to continue. From (7), let $\Psi$ denote the CDF of $h(Z)$. If $\Psi$ is continuous and strictly increasing and $\theta_{2}\in\mathbb{R}_{+}$, (29) can be expressed in closed-form as
$t^{\Theta}(x)=\theta_{1}+\theta_{2}\Psi^{-1}(F_{\mu}(x)).$ (31)
To see this, since $\theta_{2}>0$, observe that
where the last equality follows from (29) and that $F_{\nu^{\Theta}}$ is continuous and strictly increasing; here, $F_{\nu^{\Theta}}$ inherits the same properties from $\Psi$ thanks to (7). Though the continuity of $\Psi$ is not needed in Lemma 3 and its Kantorovich equivalence [24, Theorem 2.18], without it equality (a) will become an inequality which harms finding closed-form $\Theta$. Then it follows that $\frac{t^{\Theta}(x)-\theta_{1}}{\theta_{2}}=\Psi^{-1}(F_{\mu}(x))$, which yields (31). On the other hand, if $\theta_{2}=0$, since $G_{\Theta}(Z)\equiv\theta_{1}\in\mathbb{R}$, we have $F_{\nu^{\Theta}}(y)=1_{\{y\geq\theta_{1}\}}$. Plugging this into (29) directly gives $t^{\Theta}(x)=\theta_{1}$ for all $x\in\mathbb{R}$. This particularly shows that (31) is also satisfied for the case $\theta_{2}=0$. Our target sub-problem becomes
@@
As will be explained below, the other sub-problem can be solved in the same way.
@@
Together with (7), (34) becomes the constrained optimization problem
$\hbox{subject to}\quad g(\theta_{1},\theta_{2}):=-\theta_{2}\leq 0.$ (35)
The corresponding first-order Karush-Kuhn-Tucker (KKT) condition is
@@
$\lambda g(\theta_{1},\theta_{2})$ $=0,$ (37)
where $\lambda\geq 0$ is the Lagrange multiplier. By direct calculation, (36) becomes
$\theta_{1}\mathbb{E}\left[\Psi^{-1}(F_{\mu}(X))\right]+\theta_{2}\mathbb{E}\left[(\Psi^{-1}(F_{\mu}(X)))^{2}\right]$
$-\mathbb{E}\left[X\Psi^{-1}(F_{\mu}(X))\right]=\lambda/2.$
Note that $F_{\mu}(X)\sim\operatorname{Uniform}[0,1]$ from [25, Lemma 1], so that the CDF of $\Psi^{-1}(F_{\mu}(X))$ is simply $\Psi$. In other words, $\Psi^{-1}(F_{\mu}(X))$ and $h(Z)$ have identical distribution. The formulas above thus simplify to
$\theta_{1}+\theta_{2}\mathbb{E}\left[h(Z)\right]-\mathbb{E}[X]=0,$
$\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))\right)\geq 0$ (39)
in (38) since $(x-x^{\prime})(\Psi^{-1}(F_{\mu}(x))-\Psi^{-1}(F_{\mu}(x^{\prime})))\geq 0$ for all $x,x^{\prime}\in\mathbb{R}$. Specifically, by taking an independent but same distribution copy $X^{\prime}$ of $X$,
$0$ $\leq\mathbb{E}\left[(X-X^{\prime})\left(\Psi^{-1}(F_{\mu}(X))-\Psi^{-1}(F_{\mu}(X^{\prime}))\right)\right]$ (40)
$=2\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))\right).$
We therefore separate the proof into two cases from (39).
Case II: $\operatorname{Cov}(X,\Psi^{-1}(F_{\mu}(X)))=0$. As (38) entails $\lambda/2=\theta_{2}\text{Var}(h(Z))$, we have $\lambda=\theta_{2}=0$ (by recalling that either $\lambda=0$ or $\theta_{2}=0$). That is, the KKT condition gives a unique candidate optimizer $(\theta_{1}^{*},\theta_{2}^{*},\lambda^{*})=(\mathbb{E}[X],0,0)$. Since $H_{J}+\lambda H_{g}$ is again given by (IV-C), which is positive definite, we conclude that $(\theta_{1}^{*},\theta_{2}^{*})=(\mathbb{E}[X],0)$ is the unique minimizer of (35).
As a remark, another way to prove is to show $H_{J}$ as (IV-C) is positive definite in both cases first, and then $J(\theta_{1},\theta_{2})$ is convex. Since both the proof need to check the Hessian matrix, there is no simpler way to prove.

### v2

@@ III Main Results
In Sec. III-A, we analytically solve WGAN (1) with the non-linear generator and non-Gaussian data under $d=1$ when $q=2$. With large $d$ in Sec. III-B, we show that the linear generators can be asymptotically optimal for both original and unprojected sliced WGANs (6) and (7) with non-Gaussian data by solving corresponding upper bounds. We summarize and compare our results with the prior works [12, 29, 3] in Table I.
III-A Basic results for classical WGAN under one dimension
@@
We identify properties of data PDF (9) such that the optimal parameters $\theta_{2}$ have different forms (10) and (11). The proof for proposition 1 is similar to that for the upcoming Theorem (1) and given in Appendix A.
@@
where $\Phi$ is the CDF of Gaussian $N(0,1)$. However, now $\Psi$ has a jump from 0 to $1/2$ at $v=0$ and does not meet the setting of Proposition 1 since it is neither continuous nor strictly increasing. We need the following modification.
@@
Assume $\mu$ has the setting as in Proposition 1 and $\Psi$ is given by (12). If
$\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)\geq\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right),$ (13)
@@
Since the CDF of the ReLU function has a jump, the transfer function will be a piecewise function depending on the CDF in each subproblem. By solving Karush–Kuhn–Tucker (KKT) conditions of each sub-problem, the solution of the first sub-problem is indeed (14) while that of the second sub-problem is (15). We derive the closed-form primal optimal solution by eliminating the Lagrange multipliers, thereby reducing the KKT equations to a set of conditions involving only the primal variables. The condition (13) is obtained by comparing the values of the two subproblems. The full proof for Theorem 1 is in Sec. IV-A.
For the linear generator $h(Z)=Z$ as [30, eqn. (27)][12][29], one can simplify the results in Proposition 1 and also get alternative closed-form formula as follows.
@@
Assume $\mu$ has the setting as in Proposition 1 and consider the linear case $h(Z)=Z$ in (8). Then population WGAN (1) has a unique minimizer for $(\theta_{1},\theta_{2})\in\mathbb{R}\times\mathbb{R}$ as
@@
The proof of Corollary 1 is in Appendix C.
Here we briefly compare our proposed WGAN solutions with the results for the LQG setting in [12, Theorem 1]. Unlike [12], our results can deal with non-Gaussian data distribution. Moreover, we can recover the result in [12] when $q=2,d=1$. Indeed, if $X\sim\mathcal{N}(0,\sigma^{2})$, by looking into (17), we have $\Phi^{-1}(F_{\mu}(X))=\Phi^{-1}(\Phi(X/\sigma))=X/\sigma.$ Plugging this into our solution in (17) shows that $\theta_{1}^{*}=0$ and $\theta_{2}^{*}=\mathbb{E}[X\frac{X}{\sigma}]=\sigma$, coinciding with the result in [12] for $d=r=1$. We emphasize that neither our results nor those in [12] subsume the other as a special case, as [12] considers general dimension $d$ while our work is not restricted to Gaussian data distribution. Since WGAN output is already Gaussian $\Theta Z\sim\mathcal{N}(0,\mathbf{K}^{\prime})$, it is shown in [12, Theorem 1] that for this benchmark, the optimal WGAN solution happens to be the $r$-PCA solution of Gaussian $X$. That is, the optimal generator matrix $\Theta$ fulfills that $\mathbf{K}^{\prime}=\Theta\Theta^{T}$ is a rank $r$ matrix and $\mathbf{K}^{\prime}$ and $\mathbf{K}$ share the same largest $r$ eigenvalues and the corresponding eigenvectors.
Finally, we present our results for non-linear generators and first-order Wasserstein distance ($q=1$). We have
@@
Following the settings in Proposition 1, the minimizer $(\theta^{*}_{1},\theta^{*}_{2})$ of the population WGAN (1) with $q=1,d=1$ meet the following necessary conditions
@@
III-B Asymptotic optimal results for sliced WGANs under high dimension
We first focus on the asymptotic optimal generator of (6) as defined in Sec. II. Our main result shows that the linear generator can be asymptotically optimal for the original sliced WGAN (6) when $q=2$.
@@ III-B Asymptotic optimal results for sliced WGANs under high dimension
For the second-order sliced WGAN (6), if data $X$ satisfy the weak dependence condition in [25, Definition 1], the linear generator (3) is asymptotically optimal with the parameters
@@
For the proof sketch, rather than finding the closed-form solution of the original sliced WGAN (6) directly, we solve its upper bounds with bounded-error and approximations to overcome the complicated integration and projection of (6). First, by replacing general generator $G_{\Theta}(Z)$ with linear generator (3), (6) is upper-bounded by
$\min_{\Theta}\left(\int_{\omega\in\Omega}\inf_{\pi_{\omega}\in\Pi_{\omega}(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi_{\omega}}[|\omega^{T}X-\omega^{T}\Theta Z|^{q}]d\omega\right)^{\frac{1}{q}}.$ (21)
Motivated by [28], we replace the uniformly distributed $\omega\in\Omega$ in (21) with Gaussian projections $\omega\in\Omega_{G}$, that is, $\omega$ is generated according to a zero-mean Gaussian distribution with covariance matrix $(1/d)\mathbf{I}_{d}$ and aim at solving
$\min_{\Theta}\left(\int_{\omega\in\Omega_{G}}\inf_{\pi_{\omega}\in\Pi_{\omega}(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi_{\omega}}[|\omega^{T}X-\omega^{T}\Theta Z|^{q}]d\omega\right)^{\frac{1}{q}},$ (22)
@@
$\min_{\Theta}\left(\int_{\omega\in\Omega_{G}}\inf_{\pi\in\Pi(\mathcal{N}(0,\tilde{\sigma}^{2}_{x}),\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\hat{X}_{G}-\omega^{T}\Theta Z|^{q}]d\omega\right)^{1/q}$ (23)
is bounded by a function $f^{\mu}(d)=O(d^{-1/8})$, where the marginals of $\pi$ are $\mathcal{N}(0,\tilde{\sigma}^{2}_{x})$ form $\hat{X}_{G}$ and $\nu^{\Theta}_{\omega}$ from $\omega^{T}\Theta Z$. Solving (23) with the linear generator (3), the optimal $\Theta$ equals to that of
@@
As summary to the current steps, the upper-bounds of (6) under $q=2$ are
$\min_{\Theta}\left(\int_{\omega\in\Omega}\inf_{\pi_{\omega}\in\Pi_{\omega}(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi_{\omega}}[|\omega^{T}X-\omega^{T}G_{\Theta}(Z)|^{2}]d\omega\right)^{\frac{1}{2}}$
$\leq$ $\min_{\Theta}\left(\int_{\omega\in\Omega_{G}}\inf_{\pi_{\omega}\in\Pi_{\omega}(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi_{\omega}}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}]d\omega\right)^{\frac{1}{2}}$
@@
where the optimal $\Theta$ in (III-B) can be solved by (25). Then, let $r=d$, the final upper-bound in (III-B) becomes
@@
Indeed one can show that (27) tends to zero as $d\rightarrow\infty$, and thus linear generator in the second upper-bound of (III-B) is asymptotically optimal for (6). Thus, in the sense of asymptotic optimality defined in Sec. II, even with the non-linear generator, the value of sliced WGAN under very high-dimensional non-Gaussian data cannot be smaller than that from the proposed linear generator (20).
@@
Then, we prove that the objective of (28) is Schur-convex [22], so the global minimum of (28) is achieved when all the diagonal elements of $\mathbf{S}$ are equal. Thus, we find the optimal parameters (20) of the linear generator, and obtain the upper bound value (27). The full proof of solving the upper bound is given in Sec. IV-B.
@@
From [25], under the weak dependence condition, the elements of the data vector $X$ can follow a Bernoulli shift, an autoregressive (AR) model, and a Gaussian process. For channel prediction, the AR model-based method is widely used by using past channel data without requiring current channel measurements via pilot [36].
@@
As a sanity check, when $d=2$ and $q=2$, (23) with linear generator can also be solved by elliptic integrals. The solution is the same as (20). The proof is given in appendix E.
@@
Under the assumption of Theorem 2 with $\mu$ vanishing on hypersurfaces in $\mathbb{R}^{d}$, for the unprojected sliced WGAN (7), the linear generator (3) with the same parameters as (20) is asymptotically optimal when $q=2$.
@@
We show that (7) is indeed equivalent to (6) when $q=2$ by the following lemma.
@@
For any $\omega\in\Omega$, the values of the inner-discriminator probelms of (7) and (6) with the linear generator are the same when $q=2$,
$\inf_{\pi^{\prime}\in\Pi^{\prime}(\mu,\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi^{\prime}}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}]=\inf_{\pi_{\omega}\in\Pi_{\omega}(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi_{\omega}}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}].$ (29)
Though the optimal transport functions for the L.H.S. and R.H.S. are clearly different by definition, we prove that their corresponding values are the same. First, the L.H.S. in (29) belongs to the general multi-to-one-dimensional optimal transport problems with closed-form solutions identified in [9]. Then, we identify its analytical properties to prove that the optimal transport function of the L.H.S. of (29) is well-defined, and its value equals that of the R.H.S. The detailed proof of Lemma 2 is in Sec. IV-C. Finally, we can prove the desired result for (7) following the steps to prove that for the original sliced WGAN (6) in Theorem 2. ∎
For application, our upcoming simulations in Sec. V show that the linear generator (20) can be a low-complexity generator with empirical data. For data without the weak dependence condition, the state-of-the-art WGAN solver can be obtained via gradient-flow of stochastic differential equation [18][17], and our approach can be a low-complexity initialization and has the potential for faster convergence.
@@ IV-A Proof for Theorem 1
In the following, we first solve the first sub-problem with $\theta_{2}\in\mathbb{R}_{+}$ in (16). With $d=1$, we recall the following result for the inner-discriminator problem (2) of (1), by taking $\nu^{\Theta}:=P_{G_{\Theta}(z)}$, the measure generated by the generator NN (with parameter $\Theta:=(\theta_{1},\theta_{2})$ in (8)). Let $F_{\mu}$ and $F_{\nu^{\Theta}}$ denote the CDFs of the measures $\mu$ and $\nu^{\Theta}$ on $\mathbb{R}$.
@@
For the inner discriminator problem (2) with $q=2,d=1$, $\mathbb{E}_{\mu}[P(X,\theta_{1},\theta_{2})]$ for given $(\mu,\theta_{1},\theta_{2})$ equals to (31), where $t^{\Theta}(x)$ is defined as in (30).
Now we need to find a closed-form $t^{\Theta}(x)$ to continue. From (8), let $\Psi$ denote the CDF of $h(Z)$. As $\Psi$ is neither continuous nor strictly increasing, $t^{\Theta}(x)=\theta_{1}+\theta_{2}\Psi^{-1}(F_{\mu}(x))$ does not hold in general, and we cannot directly apply the proof of Proposition 1 in Appendix A. Assume $\theta_{2}\in\mathbb{R}_{+}$, our target sub-problem becomes
@@
Under the condition $\theta_{2}\geq 0$, observe that
$\Psi\bigg(\frac{t^{\Theta}(x)-\theta_{1}}{\theta_{2}}\bigg)=\mathbb{P}\bigg(h(Z)\leq\frac{t^{\Theta}(x)-\theta_{1}}{\theta_{2}}\bigg)=\mathbb{P}\left(G_{\Theta}(Z)\leq t^{\Theta}(x)\right)=F_{\nu^{\Theta}}\left(t^{\Theta}(x)\right).$ (33)
Only for $F_{\mu}(x)>1/2$, (33) equals to $F_{\mu}(x)$ from (30). From (12),
@@
$\min_{\theta_{1}\in\mathbb{R},\theta_{2}\in\mathbb{R}_{+}}\mathbb{E}_{\mu}\left[\left|X-t^{\Theta}(X)\right|^{2}\right].$ (35)
Together with (8), the sub-problem (35) becomes the constrained optimization problem
$\begin{split}&\min_{\theta_{1},\theta_{2}\in\mathbb{R}}J(\theta_{1},\theta_{2})\\ :=&\int_{\{F_{\mu}(x)\leq 1/2\}}\Big(\theta_{1}-x\Big)^{2}F_{\mu}^{\prime}(x)dx+\int_{\{F_{\mu}(x)>1/2\}}\Big(\theta_{1}+\theta_{2}\Phi^{-1}(F_{\mu}(x))-x\Big)^{2}F_{\mu}^{\prime}(x)dx\\ &\hbox{subject to}\quad g(\theta_{1},\theta_{2}):=-\theta_{2}\leq 0.\end{split}$ (36)
We will show that even with a jump, the KKT condition of (36) can be simplified as
where $\lambda\geq 0$ is the Lagrange multiplier.
The corresponding first-order KKT conditions for (36) are
@@
$\lambda g(\theta_{1},\theta_{2})$ $=0.$ (39)
By direct calculation, (38) becomes
$\int_{\{F_{\mu}(x)\leq 1/2\}}\Big(\theta_{1}-x\Big)F_{\mu}^{\prime}(x)dx+\int_{\{F_{\mu}(x)>1/2\}}\Big(\theta_{1}+\theta_{2}\Phi^{-1}(F_{\mu}(x))-x\Big)F_{\mu}^{\prime}(x)dx=0,$ (40)
$\int_{\{F_{\mu}(x)>1/2\}}\Big(\theta_{1}+\theta_{2}\Phi^{-1}(F_{\mu}(x))-x\Big)\Phi^{-1}(F_{\mu}(x))F_{\mu}^{\prime}(x)dx=\frac{\lambda}{2}.$ (41)
Recall that $X$ is a random variable with CDF $F_{\mu}$. Then, (40) and (41) can be written as
and
$\theta_{1}\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]+\theta_{2}\mathbb{E}\left[(\Phi^{-1}(F_{\mu}(X)))^{2}1_{\{F_{\mu}(X)>1/2\}}\right]$ (43)
$-\mathbb{E}\left[X\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]-\lambda/2=0,$
respectively. Plugging
$\theta_{1}=\mathbb{E}[X]-\theta_{2}\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right],$
(from (42)), into (43), we obtain
$\mathbb{E}[X]\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]+\theta_{2}\text{Var}\left(\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)$
$-\mathbb{E}\left[X\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]-\lambda/2=0.$ (44)
Hence, a solution $(\theta_{1},\theta_{2},\lambda)$ to (38) must equivalently satisfy (42) and (44). Note that $F_{\mu}(X)\sim\operatorname{Uniform}[0,1]$ from [31, Lemma 1], so that the CDF of $\Phi^{-1}(F_{\mu}(X))$ is simply $\Phi$. That is,
$\Phi^{-1}(F_{\mu}(X))\overset{(d)}{=}Z\sim\mathcal{N}(0,1),$ (45)
where $\overset{(d)}{=}$ denotes equality in distribution. It follows that
$\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]=\mathbb{E}[Z1_{\{\Phi(Z)>1/2\}}]=\mathbb{E}[Z1_{\{Z>0\}}]=\frac{1}{\sqrt{2\pi}},$ (46)
$\mathbb{E}\left[(\Phi^{-1}(F_{\mu}(X)))^{2}1_{\{F_{\mu}(X)>1/2\}}\right]=\mathbb{E}[Z^{2}1_{\{\Phi(Z)>1/2\}}]=\mathbb{E}[Z^{2}1_{\{Z>0\}}]=\frac{1}{2}.$ (47)
In view of this, (42) and (44) reduce to (37).